Back to all reports

At Junction of Queensway and High St Nth

Reported via mobile in the Traffic Lights - Permanent category anonymously at 07:48, Saturday 8 May 2021

Sent to Central Bedfordshire Council less than a minute later

The new traffic lights are still not working and the temp lights are not syncd properly and causing major congestion. Hurry up and get the new lights switched on

RSS feed of updates to this problem Receive email when updates are left on this problem.

Only the original reporter may leave updates. If you made the original report please log in to leave an update.